OpenStudy (anonymous):

integerate cos^(-1)dx

OpenStudy (anonymous):

let I = ∫cos^-1(x)dx let u = cos^-1(x), then du = -1/sqrt(1-x^2) let dv = dx, then v = x and substitute into the integration by parts formula, it should simplify down to: I = xcos^-1(x) + ∫x/sqrt(1-x^2)dx now let u = 1-x^2, then du = -2xdx, and xdx = -1/2du I = xcos^-1(x) -1/2 ∫du/sqrt(u) I = xcos^-1(x) -sqrt(u) + c I = xcos^-1(x) -sqrt(1-x^2) + c

OpenStudy (anonymous):

thanks

hartnn (hartnn):

@╰☆╮Openstudier╰☆╮ whats the source ??

OpenStudy (sidsiddhartha):

copied from yahoo lol @@╰☆╮Openstudier╰☆╮

OpenStudy (anonymous):

Yes i forgot to mention i took it from yahoo answers I know only differentiatiom not integration

hartnn (hartnn):

its always good to cite the source, else its called plagiarism! which is against the code of conduct.

OpenStudy (sidsiddhartha):

u dont need to know integration from 10th grade for IIT @╰☆╮Openstudier╰☆╮

OpenStudy (anonymous):

We r learning it tommorow!

OpenStudy (sidsiddhartha):

by the way u are from ?@╰☆╮Openstudier╰☆╮

OpenStudy (anonymous):

I am from mumbai u?

OpenStudy (sidsiddhartha):

live in kharagpur but from kolkata

OpenStudy (anonymous):

u r an iitian

OpenStudy (sidsiddhartha):

studying in 3rd year

hartnn (hartnn):

b. tech in ... ?

OpenStudy (anonymous):

i think hartn is too from kharagpur

OpenStudy (anonymous):

ya b.tech in

OpenStudy (sidsiddhartha):

EEE(electrical and electronics )